OpenStudy (anonymous):

What are the ordered pairs 2x+7y=4 12y=8-4x

OpenStudy (cwtan):

2x=4-7=-3 x=-3/2 12y=8-4(-3/2) 12y=8+6 y=7/6

OpenStudy (anonymous):

I wrote the first equation without the y in it. sorry

OpenStudy (ajprincess):

 Multiply the first equation by 2. Then subtract 2nd equation frm it.